Solve the given initial value problem for y = f(x). dy = 5x - 3 where y = -3 when x = -7. dx y
Answer #1

dy du 5* - 3 al 20 (sondes 3 y=85x-3) dx at y = 57 - 3x+ c (x=-8, y=-3 5 (12) 3 (-2) + c c = - 293 - 292 :- 9= Sx² sx 3x 293